HETAV PATEL

Grade 11,

prove that (1²×2²÷1!)+(2²×3²÷3!)+--- = 27e

prove that (1²×2²÷1!)+(2²×3²÷3!)+--- = 27e

Grade:12

3 Answers

Mohammed Anas
9 years ago
I think that is wrong...
 
And plz continue the equation for a little more next time 
 
Thanks
Mohammed 
Askiitans Student
vamshiakena
9 years ago
n^4/n!+n^2/n!+2(n^3/n!) =15e+2e+2(5e)=27e
vamshiakena
9 years ago
n^4/n!+n^2/n!+2(n^3/n!) =15e+2e+2(5e)=27e